The aeguill package

Denis Roegel

                  roegel@loria.fr

22 May 2000

When using the T1 encoding and the ae fonts, one is faced with the problem that the ae fonts do not contain the guillemets, which are necessary for the French.

This package proposes a solution to this problem. Example of french guillemets: � test �. The guillemets in the previous example are taken from the plr fonts (Polish

CMR), which are similar to EC guillemets and of which there is a Type 1 version. Hence, you will get a nice output with pdflatex.

There are four options to the aeguill package:

• ‘pl’ (default): with this, the guillemets are taken from the plr fonts.

• ‘cyr’: with this, the guillemets are taken from the wncyr fonts.

• ‘cm’: with this, the guillemets are built with the lasy fonts.

• ‘ec’: with this, the guillemets are those in the ec fonts.

The three first solutions have free Type 1 equivalents, the last doesn’t. Caution:

• If the babel package is used with the french option, do not use \CyrillicGuillemets.

• If the french package is used, you may write “<<” and “>>” instead of “«” and “»” (i.e., use 7-bit encoding instead of 8-bit encoding).